Jump to main content.


Drake Chemical

EPA Daily Situation Reports for February 1999

U. S. ARMY CORPS OF ENGINEERS
DRAKE CHEMICAL SUPERFUND SITE

182 MYRTLE STREET
LOCK HAVEN, PA 17745
PHONE 717-748-4505
FAX 717-748-7549

Sunday, February 28, 1999

FOR DATES: From 7:00 a.m. Saturday, Feb. 27, 1999 to 7:00 a.m. Sunday, Feb. 28, 1999.

WEATHER: Overcast with intermittent rain showers. The temperature range was 19oF to 42oF.

I. CURRENT SITUATION: OHM is working 24 hours a day, seven days a week operating and maintaining the Thermal Destruction Facility (TDF). The TDF was on-line for 23.9 hours and 583.9 tons of contaminated soils were thermally treated during the time frame of this report. The TDF processed contaminated soil at an average rate of 24.4 tons per hour while on line. Related ongoing site activities include feed preparation and debris separation.

II. ACTIONS TAKEN: The contractor continued TDF operations, Waste Water Treatment Plant (WWTP) operations, feed preparation, debris handling, debris separation, lime mixing, wood debris management, 2' bottom rock placement, and ash handling. Excavation was performed between stations 7+00 and 7+25 from the Con-Rail side and towards the center of the site. Continued with the real time perimeter air monitoring (PAM) program.

Received bottom ash analytical reports (bot-ash-2-18-99-368, bot-ash-2-19-99-369, bot-ash-2-20-99-370). Analytical test results indicated that thermally treated soil met the project requirements for restricted backfilling on-site (i.e. above elevation 553) due to the laboratory concentrations of arsenic, mercury, selenium, and chromium indicating less than 25 times the PADEP drinking water standard but not less than one times the standard.

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, excavation, drum handling (if encountered), feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, rock cleaning & inorganic debris decontamination and implementation of site close out plan.

IV. PROBLEMS NOTED OR ANTICIPATED: None.

V. NOTED VISITORS: None.

PREPARED BY: Richard Lupold for William Werntges (USACE)


Saturday, February 27, 1999

FOR DATES: From 7:00 a.m. Friday, Feb. 26, 1999 to 7:00 a.m. Saturday, Feb. 27, 1999.

WEATHER: Sunny and clear. The temperature range was 20oF to 45oF.

I. CURRENT SITUATION: OHM is working 24 hours a day, seven days a week operating and maintaining the Thermal Destruction Facility (TDF). The TDF was on-line for 23.9 hours and 580.9 tons of contaminated soils were thermally treated during the time frame of this report. The TDF processed contaminated soil at an average rate of 24.3 tons per hour while on line. Related ongoing site activities include feed preparation and debris separation.

II. ACTIONS TAKEN: The contractor continued TDF operations, Waste Water Treatment Plant (WWTP) operations, feed preparation, heavy equipment maintenance, equipment demobilization, debris handling, debris separation, lime mixing, wood debris management, shot rock backfilling and ash handling. A survey of the excavation bottom was performed. Minor excavation was performed along the southern face of the exclusion zone near the center of the site. Continued with the real time perimeter air monitoring (PAM) program.

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, excavation, drum handling (if encountered), feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, rock cleaning & inorganic debris decontamination and implementation of site close out plan.

IV. PROBLEMS NOTED OR ANTICIPATED: None.

V. NOTED VISITORS: None.

PREPARED BY: Steve Gibson for William Werntges (USACE)


Friday, February 26, 1999

FOR DATES: From 7:00 a.m. Thursday, Feb. 25, 1999 to 7:00 a.m. Friday, Feb. 26, 1999.

WEATHER: Partly cloudy, light snow. The temperature range was 22oF to 33oF.

I. CURRENT SITUATION: OHM is working 24 hours a day, seven days a week operating and maintaining the Thermal Destruction Facility (TDF). The TDF was on-line for 23.9 hours and 585.2 tons of contaminated soils were thermally treated during the time frame of this report. The TDF processed contaminated soil at an average rate of 24.5 tons per hour while on line. Related ongoing site activities include feed preparation and debris separation.

II. ACTIONS TAKEN: The contractor continued TDF operations, Waste Water Treatment Plant (WWTP) operations, feed preparation, wood debris management, site housekeeping, debris handling, debris separation, lime mixing, ash grading and ash handling. The contractor continued to implement the wastewater management close out plan. Excavation was performed directly south of the lime silo. The weekly NPDES water sample was collected in the WWTP. Continued with the real time perimeter air monitoring (PAM) program.

Received fly ash analytical report (fly-ash-02-17-99-073) for the contents of bin A. Analytical test results indicated that the fly ash materials met the project requirements for restricted backfilling on-site (i.e. greater than elevation 553) due to the concentration of chromium and selenium being greater than the PADEP drinking water standard, but less than 25 times the standard.

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, excavation, drum handling (if encountered), feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, rock cleaning & inorganic debris decontamination and the implementation of the project close out.

IV. PROBLEMS NOTED OR ANTICIPATED: None.

V. NOTED VISITORS: None.

PREPARED BY: Steve Gibson for Chad Thompson (USACE)


Thursday, February 25, 1999

FOR DATES: From 7:00 a.m. Wednesday, Feb. 24, 1999 to 7:00 a.m. Thursday, Feb. 25, 1999.

WEATHER: Sunny with scattered clouds. The temperature range was 17oF to 36oF.

I. CURRENT SITUATION: OHM is working 24 hours a day, seven days a week operating and maintaining the Thermal Destruction Facility (TDF). The TDF was on-line for 23.8 hours and 582.1 tons of contaminated soils were thermally treated during the time frame of this report. The TDF processed contaminated soil at an average rate of 24.5 tons per hour while on line. Related ongoing site activities include feed preparation and debris separation.

II. ACTIONS TAKEN: The contractor continued TDF operations, Waste Water Treatment Plant (WWTP) operations, feed preparation, debris handling, debris separation, lime mixing, rock cleaning, rock backfilling, ash handling and ash backfilling. A survey for the frac tank's rock foundation was performed on the 2-foot rock bottom. The contractor continued to implement the wastewater management close out plan. No new excavation was performed on this date. One roll-off containing non-hazardous materials (PPE and miscellaneous metal) was shipped off-site to a hazardous waste landfill. Continued with the real time perimeter air monitoring (PAM) program and began the weekly OSHA method 93 24-hour air sampling event for BNA at the on-site and off-site air monitoring stations.

Received bottom ash analytical report (bot-ash-02-17-99-367) for the contents of bin 7. Analytical test results indicated that the thermally treated soil met the project specifications for unrestricted backfilling on-site (i.e. above elevation 547).

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, excavation, drum handling (if encountered), feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, rock cleaning & inorganic debris decontamination and the implementation of the project close out.

IV. PROBLEMS NOTED OR ANTICIPATED: None.

V. NOTED VISITORS: Dave Mummert (OHM/IT CIH).

PREPARED BY: Steve Gibson for Chad Thompson (USACE)


Wednesday, February 24, 1999

FOR DATES: From 7:00 a.m. Tuesday, Feb. 23, 1999 to 7:00 a.m. Wednesday, Feb. 24, 1999.

WEATHER: Partly cloudy and cold. The temperature range was 8oF to 29oF.

I. CURRENT SITUATION: OHM is working 24 hours a day, seven days a week operating and maintaining the Thermal Destruction Facility (TDF). The TDF was on-line for 22.4 hours and 552.8 tons of contaminated soils were thermally treated during the time frame of this report. Off-line time to the incinerator was primarily attributed to the stoppage of feed in order to replace several sheared bolts on the quench tower screw conveyor. The TDF processed contaminated soil at an average rate of 24.7 tons per hour while on line. Related ongoing site activities include feed preparation and debris separation.

II. ACTIONS TAKEN: The contractor continued TDF operations, Waste Water Treatment Plant (WWTP) operations, feed preparation, debris handling, debris separation, lime mixing, shot rock backfilling, and ash handling. The corridor sheetpiling was cut to 2-feet below existing grade for backfill area mass grading purposes. A survey of the excavation bottom was performed. The contractor continued to implement the wastewater management close out plan. Excavation activities were continued in the exclusion zone adjacent to the corner of the ash stabilization pad. Continued with the real time perimeter air monitoring (PAM) program.

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, excavation, drum handling (if encountered), feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock cleaning & inorganic debris decontamination.

IV. PROBLEMS NOTED OR ANTICIPATED: None.

V. NOTED VISITORS: Sandra Downs (PADEP).

PREPARED BY: Steve Gibson for William Werntges (USACE)


Tuesday, February 23, 1999

FOR DATES: From 7:00 a.m. Monday, Feb. 22, 1999 to 7:00 a.m. Tuesday, Feb. 23, 1999.

WEATHER: Clear and cold. The temperature range was 9oF to 29oF.

I. CURRENT SITUATION: OHM is working 24 hours a day, seven days a week operating and maintaining the Thermal Destruction Facility (TDF). The TDF was on-line for 23.8 hours and 575.0 tons of contaminated soils were thermally treated during the time frame of this report. The TDF processed contaminated soil at an average rate of 24.2 tons per hour while on line. Related ongoing site activities include feed preparation and debris separation.

II. ACTIONS TAKEN: The contractor continued TDF operations, Waste Water Treatment Plant (WWTP) operations, feed preparation, debris handling, debris separation, lime mixing, rock cleaning, rock backfilling, ash handling, and ash backfilling. A survey of the 2-foot rock bottom and excavation bottom was performed. Excavation activities were continued in the exclusion zone adjacent to the ash stabilization pad. Continued with the real time perimeter air monitoring (PAM) program.

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, excavation, drum handling (if encountered), feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock cleaning & inorganic debris decontamination.

IV. PROBLEMS NOTED OR ANTICIPATED: None.

V. NOTED VISITORS: None.

PREPARED BY: Steve Gibson for William Werntges (USACE)


Monday, February 22, 1999

FOR DATES: From 7:00 a.m. Sunday, Feb. 21, 1999 to 7:00 a.m. Monday, Feb. 22, 1999.

WEATHER: Overcast and cold. The temperature range was 13oF to 32oF.

I. CURRENT SITUATION: OHM is working 24 hours a day, seven days a week operating and maintaining the Thermal Destruction Facility (TDF). The TDF was on-line for 23.95 hours and 598.8 tons of contaminated soils were thermally treated during the time frame of this report. The TDF processed contaminated soil at an average rate of 25 tons per hour while on line. Related ongoing site activities include feed preparation and debris separation.

II. ACTIONS TAKEN: The contractor continued TDF operations, excavation, feed preparation, equipment decontamination, debris handling, debris separation, lime mixing, rock cleaning, rock backfilling, ash handling, and ash backfilling. Continued with the real time perimeter air monitoring (PAM) program. No analytical test results for bottom ash or fly ash were received during this reports coverage.

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, excavation, drum handling (if encountered), feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock cleaning & inorganic debris decontamination.

IV. PROBLEMS NOTED OR ANTICIPATED: None.

V. NOTED VISITORS: None.

PREPARED BY: Danny Zion for Chad Thompson (USACE)


Sunday, February 21, 1999

FOR DATES: From 7:00 a.m. Saturday, Feb. 20, 1999 to 7:00 a.m. Sunday, Feb. 21, 1999.

WEATHER: Partly cloudy and cold. The temperature range was 21oF to 35oF.

I. CURRENT SITUATION: OHM is working 24 hours a day, seven days a week operating and maintaining the Thermal Destruction Facility (TDF). The TDF was on-line for 23.93 hours and 603.1 tons of contaminated soils were thermally treated during the time frame of this report. The TDF processed contaminated soil at an average rate of 25.2 tons per hour while on line. Related ongoing site activities include feed preparation and debris separation.

II. ACTIONS TAKEN: The contractor continued TDF operations, feed preparation, equipment decontamination, debris handling, debris separation, lime mixing, rock cleaning, rock backfilling, ash handling, and ash backfilling. Continued with the real time perimeter air monitoring (PAM) program. No analytical test results for bottom ash or fly ash were received during this reports coverage.

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, excavation, drum handling (if encountered), feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock cleaning & inorganic debris decontamination.

IV. PROBLEMS NOTED OR ANTICIPATED: None.

V. NOTED VISITORS: None.

PREPARED BY: Richard Lupold for William Werntges (USACE)


Saturday, February 20, 1999

FOR DATES: From 7:00 a.m. Friday, Feb. 19, 1999 to 7:00 a.m. Saturday, Feb. 20, 1999.

WEATHER: Partly cloudy. The temperature range was 22oF to 38oF.

I. CURRENT SITUATION: OHM is working 24 hours a day, seven days a week operating and maintaining the Thermal Destruction Facility (TDF). The TDF was on-line for 23.93 hours and 611.0 tons of contaminated soils were thermally treated during the time frame of this report. The TDF processed contaminated soil at an average rate of 25.5 tons per hour while on line. Related ongoing site activities include feed preparation and debris separation.

II. ACTIONS TAKEN: The contractor continued TDF operations, Waste Water Treatment Plant (WWTP) operations, feed preparation, equipment decontamination, debris handling, debris separation, lime mixing, rock cleaning, rock backfilling, ash handling, and ash backfilling. A survey of the excavation bottom and the 2-foot rock bottom was performed. Excavation activities were continued adjacent to the ash stabilization pad sheet steel. One roll-off containing hazardous waste (asbestos piping) and one roll-off containing non-hazardous material (PPE and miscellaneous metal) were shipped off-site to a hazardous waste landfill. Continued with the real time perimeter air monitoring (PAM) program. OHM began installing components of the wastewater management system to be used during the project close out. Received bottom ash analytical report (bot-ash-02-15-99-365) for the contents of bin 3. Analytical test results indicated that the thermally treated soil met the project requirements for unrestricted backfilling on-site (i.e. above elevation 547).

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, excavation, drum handling (if encountered), feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock cleaning & inorganic debris decontamination.

IV. PROBLEMS NOTED OR ANTICIPATED: None.

V. NOTED VISITORS: None.

PREPARED BY: Steve Gibson for William Werntges (USACE)


Saturday, February 19, 1999

FOR DATES: From 7:00 a.m. Thursday, Feb. 18, 1999 to 7:00 a.m. Friday, Feb. 19, 1999.

WEATHER: Overcast. The temperature range was 31oF to 41oF.

I. CURRENT SITUATION: OHM is working 24 hours a day, seven days a week operating and maintaining the Thermal Destruction Facility (TDF). The TDF was on-line for 23.8 hours and 651.5 tons of contaminated soils were thermally treated during the time frame of this report. The TDF processed contaminated soil at an average rate of 27.3 tons per hour while on line. Related ongoing site activities include feed preparation and debris separation.

II. ACTIONS TAKEN: The contractor continued TDF operations, Waste Water Treatment plant operations and maintenance, feed preparation, debris handling, debris separation, lime mixing, rock cleaning, rock and shot rock backfilling, ash handling and ash backfilling. Excavation activities were continued along the ash pad sheet steel and adjacent to the Conrail sheet piling. Continued with the real time perimeter air monitoring (PAM) program.

Received bottom ash analytical report (bot-ash-02-14-99-364) for the contents of bin 3 and fly ash analytical report (flyash-2-13-99-072) for the contents of bin E. Analytical test results indicated that the ash met the project requirements for restricted backfilling on-site (i.e. above elevation 553) due to the concentrations of chromium in the bottom ash and cadmium, chromium and selenium in the fly ash being greater than the PADEP drinking water standard, but less than 25 times the standard.

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, excavation, drum handling (if encountered), feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock cleaning & inorganic debris decontamination.

IV. PROBLEMS NOTED OR ANTICIPATED: None.

V. NOTED VISITORS: None.

PREPARED BY: Danny Zion for Chad Thompson (USACE)


Thursday, February 18, 1999

FOR DATES: From 7:00 a.m. Wednesday, Feb. 17, 1999 to 7:00 a.m. Thursday, Feb. 18, 1999.

WEATHER: Overcast with light rain. The temperature range was 31oF to 44oF.

I. CURRENT SITUATION: OHM is working 24 hours a day, seven days a week operating and maintaining the Thermal Destruction Facility (TDF). The TDF was on-line for 23.9 hours and 697.5 tons of contaminated soils were thermally treated during the time frame of this report. The TDF processed contaminated soil at an average rate of 29.2 tons per hour while on line. Related ongoing site activities include feed preparation and debris separation.

II. ACTIONS TAKEN: The contractor continued TDF operations, Waste Water Treatment plant operations and maintenance, feed preparation, debris handling, debris separation, lime mixing, rock cleaning, rock and shot rock backfilling, ash handling and ash backfilling. Excavation activities were continued along the ash pad sheet steel and adjacent to the Conrail sheet piling. Drum bulking activities were continued in the exclusion zone. On this date, the contents of four non-reactive drums were bulked for homogenization and sampling. Continued with the real time perimeter air monitoring (PAM) program.

Received bottom ash analytical report (bot-ash-02-13-99-363) for the contents of bin 5. Analytical test results indicated that thermally treated soil did not meet the project requirements for backfilling on-site due to the concentration of Beta-Naphthylamine being greater than the established organic criteria. The contents of bin 5 will be returned to the exclusion zone for further thermal treatment. The contractor has implemented several process changes to lessen the probability of future bottom ash bins not meeting site back fill criteria. Process changes include lowering the waste feed rate, increasing the kiln exit gas and ash discharge temperatures, and making adjustments to the kiln rotational speed.

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, excavation, drum handling (if encountered), feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock cleaning & inorganic debris decontamination.

IV. PROBLEMS NOTED OR ANTICIPATED: None.

V. NOTED VISITORS: Ross Wise (USACE)

PREPARED BY: Steve Gibson for Chad Thompson (USACE)


Wednesday, February 17, 1999

FOR DATES: From 7:00 a.m. Tuesday, Feb. 16, 1999 to 7:00 a.m. Wednesday, Feb. 17, 1999.

WEATHER: Clear and mild. The temperature range was 22oF to 56oF.

I. CURRENT SITUATION: OHM is working 24 hours a day, seven days a week operating and maintaining the Thermal Destruction Facility (TDF). The TDF was on-line for 23.9 hours and 836.2 tons of contaminated soils were thermally treated during the time frame of this report. The TDF processed contaminated soil at an average rate of 35.0 tons per hour while on line. Related ongoing site activities include feed preparation and debris separation.

II. ACTIONS TAKEN: The contractor continued TDF operations, Waste Water Treatment plant operations and maintenance, feed preparation, debris handling, debris separation, lime mixing, rock cleaning, rock backfilling, ash handling and ash backfilling. A confirmation survey of the excavation bottom and the 2-foot rock bottom was performed. The activated carbon, from two of the carbon vessels used in the WWTP, was exchanged with new activated carbon. The old carbon was shipped off-site for recycling. Excavation activities were continued along the ash pad sheet steel. Continued with the real time perimeter air monitoring (PAM) program.

Received bottom ash analytical report (bot-ash-02-13-99-362) for the contents of bin 6. Analytical test results indicated that thermally treated soil met the project requirements for unrestricted backfilling on-site (i.e. above elevation 547).

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, excavation, drum handling (if encountered), feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock cleaning & inorganic debris decontamination.

IV. PROBLEMS NOTED OR ANTICIPATED: None.

V. NOTED VISITORS: Sandra Downs and Larry Newcomer (PADEP)

PREPARED BY: Steve Gibson for William Werntges (USACE)


Tuesday, February 16, 1999

FOR DATES: From 7:00 a.m. Monday, Feb. 15, 1999 to 7:00 a.m. Tuesday, Feb. 16, 1999.

WEATHER: Clear. The temperature range was 19oF to 50oF.

I. CURRENT SITUATION: OHM is working 24 hours a day, seven days a week operating and maintaining the Thermal Destruction Facility (TDF). The TDF was on-line for 23.8 hours and 755.9 tons of contaminated soils were thermally treated during the time frame of this report. The TDF processed contaminated soil at an average rate of 31.8 tons per hour while on line. Related ongoing site activities include feed preparation and debris separation.

II. ACTIONS TAKEN: The contractor continued TDF operations and maintenance, excavation, feed preparation, debris handling, debris separation, lime mixing, ash handling, ash backfilling, and WWTP operations. Continued with the real time perimeter air monitoring (PAM) program.

Received bottom ash analytical reports (bot-ash-02-10-99-360 and bot-ash-02-11-99-361) for the contents of bins 4 & 7. Analytical test results indicated that thermally treated soil met the project requirements for unrestricted backfilling on-site (i.e. above elevation 547).

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, excavation, drum handling (if encountered), feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock cleaning & inorganic debris decontamination.

IV. PROBLEMS NOTED OR ANTICIPATED: None.

V. NOTED VISITORS: Site tours were given to Williamsport Area School District high school and middle school science teachers during the morning. A site tour was given to Penn State University environmental law students during the afternoon.

PREPARED BY: Danny Zion for William Werntges (USACE)


Monday, February 15, 1999

FOR DATES: From 7:00 a.m. Sunday, Feb. 14, 1999 to 7:00 a.m. Monday, Feb. 15, 1999.

WEATHER: Clear and breezy. The temperature range was 18oF to 36oF.

I. CURRENT SITUATION: OHM is working 24 hours a day, seven days a week operating and maintaining the Thermal Destruction Facility (TDF). The TDF was on-line for 23.9 hours and 655.6 tons of contaminated soils were thermally treated during the time frame of this report. The TDF processed contaminated soil at an average rate of 27.4 tons per hour while on line. Related ongoing site activities include feed preparation and debris separation.

II. ACTIONS TAKEN: The contractor continued TDF operations and maintenance, excavation, feed preparation, debris handling, debris separation, lime mixing, ash handling, and ash backfilling. Continued with the real time perimeter air monitoring (PAM) program.

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, excavation, drum handling (if encountered), feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock cleaning & inorganic debris decontamination.

IV. PROBLEMS NOTED OR ANTICIPATED: None.

V. NOTED VISITORS: None.

PREPARED BY: Danny Zion for William Werntges (USACE)


Sunday, February 14, 1999

FOR DATES: From 7:00 a.m. Saturday, Feb. 13, 1999 to 7:00 a.m. Sunday, Feb. 14, 1999.

WEATHER: Overcast and breezy, snow in the afternoon. The temperature range was 19oF to 30oF.

I. CURRENT SITUATION: OHM is working 24 hours a day, seven days a week operating and maintaining the Thermal Destruction Facility (TDF). The TDF was on-line for 20.4 hours and 640.4 tons of contaminated soils were thermally treated during the time frame of this report. The majority of the incinerator off-line time was due to a specific gravity reading in the acid gas scrubbing system outside of the normal operating range. The scrubber water blow down rate was increased temporarily until the readings taken were within normal operating ranges, at which time waste feed to the unit was resumed. The TDF processed contaminated soil at an average rate of 31.4 tons per hour while on line. Related ongoing site activities include feed preparation and debris separation.

II. ACTIONS TAKEN: The contractor continued TDF operations and maintenance, feed preparation, debris handling, debris separation, lime mixing, ash handling, and ash backfilling. Excavation activities were continued along the southern face of the exclusion zone in the center of the site. Continued with the real time perimeter air monitoring (PAM) program and recovery of the OSHA method 93 24-hour air-sampling event for BNA.

Received bottom ash analytical report (bot-ash-2-9-99-359) for the contents of bin 3. Analytical test results indicated that the thermally treated soil met the project requirements for unrestricted backfilling on-site.

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, excavation, drum handling (if encountered), feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock cleaning & inorganic debris decontamination.

IV. PROBLEMS NOTED OR ANTICIPATED: None.

V. NOTED VISITORS: None.

PREPARED BY: Richard Lupold for William Werntges (USACE)


Saturday, February 13, 1999

FOR DATES: From 7:00 a.m. Friday, Feb. 12, 1999 to 7:00 a.m. Saturday, Feb. 13, 1999.

WEATHER: Clear and mild, freezing rain/snow in the afternoon. The temperature range was 48oF to 29oF.

I. CURRENT SITUATION: OHM is working 24 hours a day, seven days a week operating and maintaining the Thermal Destruction Facility (TDF). The TDF was on-line for 3.3 hours and 105.8 tons of contaminated soils were thermally treated during the time frame of this report. Off-line time to the incinerator is attributed to the stoppage of feed in order perform scheduled maintenance and monthly calibrations to the TDF. An additional delay to the restart of waste feed was due to a malfunction of a sensor in the emergency generator which was repaired. The TDF processed contaminated soil at an average rate of 32.0 tons per hour while on line. Related ongoing site activities include feed preparation and debris separation.

II. ACTIONS TAKEN: The contractor continued TDF operations and maintenance, feed preparation, equipment decontamination, debris handling, debris separation, lime mixing, rock cleaning, rock backfilling, ash handling, and ash backfilling. Excavation activities were continued along the southern face of the exclusion zone in the center of the site. Continued with the real time perimeter air monitoring (PAM) program and initiated the OSHA method 93 24-hour air-sampling event for BNA.

Received bottom ash analytical report (bot-ash-2-8-99-358) for the contents of bin 2. Analytical test results indicated that the thermally treated soil met the project requirements for unrestricted backfilling on-site. Received fly ash analytical report (fly-ash-2-8-99-358) for the contents of bin D. Analytical test results indicated that the fly ash materials met the project requirements for restricted backfilling on-site (i.e. above elevation 553) due to the concentrations of chromium and selenium being greater than the PADEP drinking water standard, but less than 25 times the standard.

Feed to the incinerator was shut down at 0800 hours in order to perform maintenance to the dual feed screws, perform the monthly calibrations and perform the monthly interlock checks. As part of the interlock checks, a Thermal Relief Valve (TRV) opening was initiated at 1421 hours. Hazardous waste feed to the incinerator had been off-line for approximately 6.3 hours before the TRV opening and the proper authorities were notified. Hazardous waste feed resumed at 0441 hours.

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, excavation, drum handling (if encountered), feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock cleaning & inorganic debris decontamination.

IV. PROBLEMS NOTED OR ANTICIPATED: None.

V. NOTED VISITORS: None.

PREPARED BY: Greg Beauseigneur for William Werntges (USACE)


Friday, February 12, 1999

FOR DATES: From 7:00 a.m. Thursday, Feb. 11, 1999 to 7:00 a.m. Friday, Feb.12, 1999.

WEATHER: Sunny with few clouds. The temperature range was 24oF to 47oF.

I. CURRENT SITUATION: OHM is working 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). The TDF was on-line for 23.9 hours and 832.9 tons of contaminated soils were thermally treated during the time frame of this report. The TDF processed contaminated soil at an average rate of 34.9 tons per hour while on line. Related ongoing site activities include feed preparation and debris separation.

At 0800 hours this morning, Friday, February 12, 1999, feed to the incinerator was stopped and the unit was placed in hot hold in order to perform preventative maintenance to the dual feed screws and to perform the monthly TDF calibrations. During this time period, the monthly interlock checks will be performed in order to test the TDF's automatic shut down systems. As part of the interlock checks, OHM will initiate a Thermal Relief Valve (TRV) opening during the later morning hours. Prior to the TRV opening, the unit will be purged of all hazardous waste materials and the proper authorities will be notified. Hazardous waste feed is tentatively scheduled to resume at 1700 hours.

II. ACTIONS TAKEN: The contractor continued TDF operations, Waste Water Treatment Plant (WWTP) operations, feed preparation, debris handling, debris separation, lime mixing, rock screening, rock backfilling, ash handling and ash backfilling. A survey of the excavation floor and the 2-foot rock bottom was performed. Excavation activities were continued in the exclusion zone adjacent to the ash pad sheet steel. Two roll-offs containing non-hazardous materials (PPE and dry decontaminated drum carcasses) were shipped off-site to a hazardous waste landfill. The contractor continued with the real time perimeter air monitoring (PAM) program. Received bottom ash analytical report (bot-ash-02-7-99-357) for the contents of bin 1. Analytical test results indicated that thermally treated soil met the project requirements for unrestricted backfilling on-site (i.e. above elevation 547).

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, excavation, drum handling (if encountered), feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock cleaning & inorganic debris decontamination.

IV. PROBLEMS NOTED OR ANTICIPATED: None.

V. NOTED VISITORS: Dr. Scott Tucker (IT).

PREPARED BY: Steve Gibson for Chad Thompson (USACE)


Thursday, February 11, 1999

FOR DATES: From 7:00 a.m. Wednesday, Feb. 10, 1999 to 7:00 a.m. Thursday, Feb. 11, 1999.

WEATHER: Sunny. The temperature range was 24oF to 50oF.

I. CURRENT SITUATION: OHM is working 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). The TDF was on-line for 23.9 hours and 840.5 tons of contaminated soils were thermally treated during the time frame of this report. The TDF processed contaminated soil at an average rate of 35.2 tons per hour while on line. Related ongoing site activities include feed preparation and debris separation.

II. ACTIONS TAKEN: The contractor continued TDF operations, Waste Water Treatment Plant (WWTP) operations, feed preparation, wood shredding, debris handling, debris separation, lime mixing, rock screening, rock backfilling, ash handling and ash backfilling. Excavation activities were continued in the exclusion zone adjacent to the ash pad sheet steel. The contractor continued with the real time perimeter air monitoring (PAM) program.

Received bottom ash analytical report (bot-ash-02-6-99-356) for the contents of bin 5. Analytical test results indicated that thermally treated soil met the project requirements for unrestricted backfilling on-site (i.e. above elevation 547).

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, excavation, drum handling (if encountered), feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock cleaning & inorganic debris decontamination.

IV. PROBLEMS NOTED OR ANTICIPATED: None.

V. NOTED VISITORS: Dr. Scott Tucker (IT).

PREPARED BY: Danny Zion for Chad Thompson (USACE)


Wednesday, February 10, 1999

FOR DATES: From 7:00 a.m. Tuesday, Feb. 9, 1999 to 7:00 a.m. Wednesday, Feb. 10, 1999.

WEATHER: Mostly sunny and mild. The temperature range was 24oF to 49oF.

I. CURRENT SITUATION: OHM is working 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). The TDF was on-line for 23.9 hours and 831.8 tons of contaminated soils were thermally treated during the time frame of this report. The TDF processed contaminated soil at an average rate of 34.8 tons per hour while on line. Related ongoing site activities include feed preparation and debris separation.

II. ACTIONS TAKEN: The contractor continued TDF operations, Waste Water Treatment Plant (WWTP) operations, feed preparation, wood shredding, debris handling, debris separation, lime mixing, rock screening, rock backfilling, ash handling and ash backfilling. Performed soil compaction testing in the backfill area. Excavation activities were continued in the exclusion zone adjacent to the ash pad sheet steel. The contractor continued with the real time perimeter air monitoring (PAM) program and initiated the OSHA method 93 24-hour air sampling event for BNA at the on-site and off-site air monitoring stations.

Received bottom ash analytical report (bot-ash-02-5-99-355) for the contents of bin 6. Analytical test results indicated that thermally treated soil met the project requirements for unrestricted backfilling on-site (i.e. above elevation 547).

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, excavation, drum handling (if encountered), feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock cleaning & inorganic debris decontamination.

IV. PROBLEMS NOTED OR ANTICIPATED: None.

V. NOTED VISITORS: Sandra Downs, Larry Newcomer (PADEP), Dr. Scott Tucker (IT).

PREPARED BY: Steve Gibson for Chad Thompson (USACE)


Tuesday, February 9, 1999

FOR DATES: From 7:00 a.m. Monday, Feb. 8, 1999 to 7:00 a.m. Tuesday, Feb. 9, 1999.

WEATHER: Partly cloudy to clear. The temperature range was 24oF to 37oF.

I. CURRENT SITUATION: OHM is working 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). The TDF was on-line for 23.8 hours and 842.3 tons of contaminated soils were thermally treated during the time frame of this report. The TDF processed contaminated soil at an average rate of 35.4 tons per hour while on line. Related ongoing site activities include feed preparation and debris separation.

II. ACTIONS TAKEN: The contractor continued TDF operations, Waste Water Treatment Plant (WWTP) operations, feed preparation, debris handling, debris separation, lime mixing, rock screening, ash handling and ash backfilling. A survey was performed to confirm the excavation bottom and the 2-foot rock bottom elevations. Excavation activities were continued in the exclusion zone. The contractor continued with the real time perimeter air monitoring (PAM) program.

Received bottom ash analytical reports (bot-ash-02-02-99-352B, bot-ash-02-02-99-352A and bot-ash-02-04-99-354) for the contents of bin 3, bin 4 and bin 7, respectively. Analytical test results indicated that thermally treated soil met the project requirements for unrestricted backfilling on-site (i.e. above elevation 547).

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, excavation, drum handling (if encountered), drum bulking, feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock cleaning & inorganic debris decontamination.

IV. PROBLEMS NOTED OR ANTICIPATED: None.

V. NOTED VISITORS: None.

PREPARED BY: Steve Gibson for William Werntges (USACE)


Monday, February 8, 1999

FOR DATES: From 7:00 a.m. Sunday, Feb. 7, 1999 to 7:00 a.m. Monday, Feb. 8, 1999.

WEATHER: Overcast & snow. The temperature range was 29oF to 33oF.

I. CURRENT SITUATION: OHM is working 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). The TDF was on-line for 23.9 hours and 834.4 tons of contaminated soils were thermally treated during the time frame of this report. The TDF processed contaminated soil at an average rate of 34.9 tons per hour while on line. Related ongoing site activities include feed preparation and debris separation.

II. ACTIONS TAKEN: The contractor continued TDF operations, excavation, feed preparation, debris handling, debris separation, lime mixing, rock screening, rock cleaning, rock backfilling, ash handling, and ash backfilling and drum bulking. The contractor continued with the real time perimeter air monitoring (PAM) program.

Received bottom ash analytical report (bot-ash-2-3-99-353). Analytical test results indicated that thermally treated soil met the project requirements for restricted backfilling on-site (i.e above elevation 553) due to the concentrations of lead being greater than the PADEP drinking water standard, but less than 25 times the standard.

Received fly ash analytical report (fly-ash-2-3-99-070). Analytical test results indicated that the fly ash materials met the project requirements for restricted backfilling on-site (i.e. above elevation 553) due to the concentrations of cadmium, chromium, lead, and selenium being greater than the PADEP drinking water standard, but less than 25 times the standard.

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, excavation, drum handling (if encountered), drum bulking, feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock cleaning & inorganic debris decontamination.

IV. PROBLEMS NOTED OR ANTICIPATED: None.

V. NOTED VISITORS:

PREPARED BY: Danny Zion for Chad Thompson (USACE)


Sunday, February 7, 1999

FOR DATES: From 7:00 a.m. Saturday, Feb. 6, 1999 to 7:00 a.m. Sunday, Feb. 7, 1999.

WEATHER: Overcast. The temperature range was 28oF to 46oF.

I. CURRENT SITUATION: OHM is working 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). The TDF was on-line for 23.9 hours and 841.6 tons of contaminated soils were thermally treated during the time frame of this report. The TDF processed contaminated soil at an average rate of 35.2 tons per hour while on line. Related ongoing site activities include feed preparation and debris separation.

II. ACTIONS TAKEN: The contractor continued TDF operations, feed preparation, debris handling, debris separation, lime mixing, rock screening, rock cleaning, rock backfilling, ash handling, and ash backfilling and drum bulking. Excavation activities were continued on the southern portion of the excavation face. The contractor continued with the real time perimeter air monitoring (PAM) program.

No bottom or fly ash analytical reports were received during this reports coverage.

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, excavation, drum handling (if encountered), drum bulking, feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock cleaning & inorganic debris decontamination.

IV. PROBLEMS NOTED OR ANTICIPATED: None.

V. NOTED VISITORS:

PREPARED BY: Richard Lupold for William Werntges (USACE)


Saturday, February 6, 1999

FOR DATES: From 7:00 a.m. Friday, Feb. 5, 1999 to 7:00 a.m. Saturday, Feb. 6, 1999.

WEATHER: Clear and sunny. The temperature range was 29oF to 41oF.

I. CURRENT SITUATION: OHM is working 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). The TDF was on-line for 23.9 hours and 815.7 tons of contaminated soils were thermally treated during the time frame of this report. The TDF processed contaminated soil at an average rate of 34.1 tons per hour while on line. Related ongoing site activities include feed preparation and debris separation.

II. ACTIONS TAKEN: The contractor continued TDF operations, feed preparation, debris handling, debris separation, lime mixing, rock screening, rock cleaning, rock backfilling, ash handling, and ash backfilling. Excavation activities were continued on the southern portion of the excavation face. Nestlerode Construction finished the installation of the sheet piling shoring system. The contractor continued with the real time perimeter air monitoring (PAM) program.

Received bottom ash analytical reports (bot-ash-01-31-99-350 and bot-ash-02-01-99-351) for the contents of bin 1 and bin 5, respectively. Analytical test results indicated that the thermally treated soil met the project requirements for unrestricted backfilling on-site (i.e. above elevation 547).

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, excavation, drum handling (if encountered), feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock cleaning & inorganic debris decontamination.

IV. PROBLEMS NOTED OR ANTICIPATED: None.

V. NOTED VISITORS: Sandra Downs (PADEP).

PREPARED BY: Steve Gibson for William Werntges (USACE)


Friday, February 5, 1999

FOR DATES: From 7:00 a.m. Thursday, Feb. 4, 1999 to 7:00 a.m. Friday, Feb. 5, 1999.

WEATHER: Overcast. The temperature range was 30* to 47*F.

I. CURRENT SITUATION: OHM is working 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). The TDF was on-line for 23.8 hours and 827.9 tons of contaminated soils were thermally treated during the time frame of this report. The TDF processed contaminated soil at an average rate of 34.9 tons per hour while on line. Related ongoing site activities include feed preparation and debris separation.

II. ACTIONS TAKEN: The contractor continued TDF operations, Waste Water Treatment Plant (WWTP) operations, feed preparation, debris handling, debris separation, lime mixing, rock cleaning, rock backfilling, ash handling, and ash backfilling. Excavation activities were continued directly south of the ash stabilization building in the vicinity of the former debris separation building. Nestlerode Construction continued to install the walers for the CZ113 sheet piling shoring system. The contractor continued with the real time perimeter air monitoring (PAM) program and finished the monthly air sampling event for TSP/metals and semi-volatile organic compounds at the perimeter air monitoring stations.

Received fly ash analytical report (fly-ash-01-29-99-069) for the contents of fly ash bin A. Analytical test results indicated that the fly ash materials met the project requirements for restricted backfilling on-site (i.e. above elevation 553) due to the concentrations of chromium and selenium being greater than the PADEP drinking water standard, but less than 25 times the standard.

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, excavation, drum handling (if encountered), feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock cleaning & inorganic debris decontamination.

IV. PROBLEMS NOTED OR ANTICIPATED: None.

V. NOTED VISITORS: None.

PREPARED BY: Steve Gibson for Chad Thompson (USACE)


Thursday, February 4, 1999

FOR DATES: From 7:00 a.m. Wednesday, Feb. 3, 1999 to 7:00 a.m. Thursday, Feb. 4, 1999.

WEATHER: Sunny and clear. The temperature range was 30* to 47*F.

I. CURRENT SITUATION: OHM is working 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). The TDF was on-line for 23.2 hours and 783.2 tons of contaminated soils were thermally treated during the time frame of this report. Off-line time is primarily attributed to the stoppage of feed in order to replace a worn out roller on the weighbelt conveyor. The TDF processed contaminated soil at an average rate of 33.8 tons per hour while on line. Related ongoing site activities include feed preparation and debris separation.

II. ACTIONS TAKEN: The contractor continued TDF operations, Waste Water Treatment Plant (WWTP) operations, equipment decontamination, feed preparation, debris handling, debris separation, lime mixing, rock cleaning, rock backfilling, ash handling, and ash backfilling. A survey was performed to confirm the excavation and rock bottom elevation. Four intact drums were sampled for fingerprint analyses and placed in overpack containers for temporary storage. Excavation activities were continued directly south of the ash stabilization building sheet steel. Nestlerode Construction continued to install the whalers for the CZ113 sheet piling shoring system. The contractor continued with the real time perimeter air monitoring (PAM) program and performed the monthly air sampling event for TSP/metals and semi-volatile organic compounds at the perimeter air monitoring stations.

Received bottom ash analytical report (bot-ash-01-30-349) for the contents of bin 6. Analytical test results indicated that the thermally treated soil met the project requirements for unrestricted backfilling on-site (i.e. above elevation 547).

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, excavation, drum handling (if encountered), feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock cleaning & inorganic debris decontamination. Nestlerode Construction will be on-site to install the whaler shoring system for the sheet piling in the exclusion zone.

IV. PROBLEMS NOTED OR ANTICIPATED: None.

V. NOTED VISITORS: None.

PREPARED BY: Steve Gibson for William Werntges (USACE)


Wednesday, February 3, 1999

FOR DATES: From 7:00 a.m. Tuesday, Feb. 2, 1999 to 7:00 a.m. Wednesday, Feb. 3, 1999.

WEATHER: Overcast with light rain. The temperature range was 34 to 41oF.

I. CURRENT SITUATION: OHM is working 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). The TDF was on-line for 23.9 hours and 809.1 tons of contaminated soils were thermally treated during the time frame of this report. The TDF processed contaminated soil at an average rate of 33.9 tons per hour while on line. Related ongoing site activities include feed preparation and debris separation.

II. ACTIONS TAKEN: The contractor continued TDF operations, Waste Water Treatment Plant (WWTP) operations, equipment decontamination, feed preparation, debris handling, debris separation, lime mixing, rock cleaning, rock backfilling, ash handling, and ash backfilling. Excavation activities were continued directly south of the ash stabilization building sheet steel. Nestlerode Construction continued to install the whalers for the CZ113 sheet piling shoring system. Continued with the real time perimeter air monitoring (PAM) program.

Received bottom ash analytical reports (bot-ash-01-28-99-347 and bot-ash-01-29-99-348) for the contents of bin 2 and bin 7, respectively. Analytical test results indicated that the thermally treated soil met the project requirements for unrestricted backfilling on-site (i.e. above elevation 547).

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, excavation, drum handling (if encountered), feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock cleaning & inorganic debris decontamination. Nestlerode Construction will be on-site to install the whaler shoring system for the sheet piling in the exclusion zone.

IV. PROBLEMS NOTED OR ANTICIPATED: None.

V. NOTED VISITORS: Sandra Downs (PADEP), Trina Allen (TAG), Bill Smedley (AIR)

PREPARED BY: Steve Gibson for William Werntges (USACE)


Tuesday, February 2, 1999

FOR DATES: From 7:00 a.m. Monday, Feb. 1, 1999 to 7:00 a.m. Tuesday, Feb. 2, 1999.

WEATHER: Mostly sunny with evening rain. The temperature range was 14 to33oF.

I. CURRENT SITUATION: OHM is working 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). The TDF was on-line for 23.8 hours and 838.3 tons of contaminated soils were thermally treated during the time frame of this report. The TDF processed contaminated soil at an average rate of 35.2 tons per hour while on line. Related ongoing site activities include feed preparation and debris separation.

II. ACTIONS TAKEN: The contractor continued TDF operations, Waste Water Treatment Plant (WWTP) maintenance, equipment decontamination, feed preparation, debris handling, debris separation, lime mixing, rock cleaning, rock backfilling, ash handling, and ash backfilling. Excavation activities were continued beneath the former rock wash pad. Nestlerode Construction began installing the whalers for the CZ113 sheet piling shoring system. Continued with the real time perimeter air monitoring (PAM) program.

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, excavation, drum handling (if encountered), feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock cleaning & inorganic debris decontamination. Nestlerode Construction will be on-site to install the whaler shoring system for the sheet piling in the exclusion zone.

IV. PROBLEMS NOTED OR ANTICIPATED: None.

V. NOTED VISITORS: None.

PREPARED BY: Steve Gibson for William Werntges (USACE)


Monday, February 1, 1999

FOR DATES: From 7:00 a.m. Sunday, Jan. 31, 1999 to 7:00 a.m. Monday, Feb. 01, 1999.

WEATHER: Mostly sunny. The temperature range was 14 to 34oF.

I. CURRENT SITUATION: OHM is working 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). The TDF was on-line for 23.9 hours and 830.7 tons of contaminated soils were thermally treated during the time frame of this report. The TDF processed contaminated soil at an average rate of 34.8 tons per hour while on line. Related ongoing site activities include feed preparation and debris separation.

II. ACTIONS TAKEN: The contractor continued TDF operations, excavation, feed preparation, debris handling, debris separation, lime mixing, ash handling, and ash backfilling. Continued with the real time perimeter air monitoring (PAM) program.

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, excavation, drum handling (if encountered), feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock cleaning & inorganic debris decontamination. Nestlerode Construction will be on-site to install the whaler shoring system for the sheet piling in the exclusion zone.

IV. PROBLEMS NOTED OR ANTICIPATED: None.

V. NOTED VISITORS: None.

PREPARED BY: Danny Zion for William Werntges (USACE)

Region 3 | Mid-Atlantic Cleanup | Mid-Atlantic Superfund |EPA Home | EPA Superfund Homepage


Local Navigation


Jump to main content.